# The process of polarisation as a loss of dimensionality: measuring changes in polarisation using Singular Value Decomposition of network graphs

### v1

@@ 2.3 Code
All scripts are available at https://github.com/gvdr/Sage_data.
3 Data
We apply our computational framework for polarisation to three different data sets: two from Twitter, and one consisting of simulated interaction networks.
@@ 3.1 Climate discussion in New Zealand Twitter
We obtained 12939 tweets by querying Twitter’s Academic API v2.0 for keywords related to climate change: climate, pollution, age, CO2, and carbon. We restricted our search to tweets geographically tagged as being from New Zealand, and tweets published after 2017. We then considered two time windows that corresponded to roughly equally sized networks: between 2017 and 2020, and 2020 to 2023. For each time frame, we built a network by considering each user (identified by their unique IDs) as a node, and any mention, reply, or quote tweet between two users as an edge. There were 6767 tweets in the 2017-2020 network and 6172 in the 2020-2023 network. We analyzed the two networks independently.
3.2 COP discussion in Twitter
Falkenberg et al. collected a large corpus of tweets by querying Twitter’s Academic API v2.0 for tweets mentioning ”COP2x” where x was an integer between 0 and 6 (inclusive) [17]. They restricted their search to tweets in English, and covered the COP from 20 to 26 (years 2014 to 2022, with 2020 and 2021 skipped due to the Covid-19 pandemic). Their adjacency matrix was constructed based on whether a user $i$ retweeted tweets from a political influencer $j$. Their focus was whether there was noticeable division among users based on whether they were spreading true information about climate change or disinformation from climate change denialist influencers.
To test for polarisation using our framework, we built a network for each COP using the same data by considering each user (identified by their unique IDs) as a node, and any mention, reply, or quote tweet between two users as an edge. The network for each year was analysed independently.
3.3 Synthetic Data
We wanted to explore what happens when common axioms about polarisation are evaluated using our approach. To do this, we generated two networks using stochastic block models. The first network is used to evaluate the how changing the chances of between-groups engagement affects dimensionality of the network. The second dataset is used to evaluate what happens when the two groups in a polarisation are different sizes, to test the concept that polarisation is maximised when the two groups are equal in size and then decreases as one group becomes predominant.
Between-groups engagement
We simulated a stochastic block network of 1000 nodes, split into two equally sized groups. We varied the probability of each node forming a connection within its group and varying probabilities of in-group and between-groups linking. In particular, we simulated a networks with in-group link probability of 0.3 to 0.45 with steps of 0.05. and between-group probabilities of 0.1, 0.05, and 0.01.
Predominance of a group
We simulated a stochastic block network of 1000 nodes, split into two groups. We fixed probabilities of in-group linking between 0.3 and 0.45 with steps of 0.05, and fixed the probability of between-groups linking at 0.05. We varied the sizes of the two groups, in particular, we simulated a process in which one of the two groups becomes predominant, progressively increasing its size from 0.5 of the full network to 0.2, 0.1, and finally 0.01.
4 Results
4.1 NZ Twitter Climate Change Data
Figure 1: Plot comparing the network dimensionality of NZ climate change tweets in 2017-2020 to the dimensionality in 2020-2023.
Figure 2: Plot comparing the von Neumann entropy of NZ climate change tweets in 2017-2020 to the entropy in 2020-2023.
@@ 4.1 NZ Twitter Climate Change Data
2017-2020 39.0 39.0 0.980229 0.979954
2020-2023 27.0 24.0 0.97439 0.97372
We find that the dimensionality of the NZ Twitter discussion of climate change has decreased over time, indicating that it is becoming more polarised. Similarly, the von Neumann entropy of the network is lower in the 2020-2023 period than in the 2017-2020 period. This suggests that the complexity of the network is decreasing and the political positions that it is possible for users to hold are becoming narrower over time, which matches Baldassarri and Gelman’s view of how polarisation works politically.
4.2 COP Data
@@
Figure 3: Plot comparing the network dimensionality of tweets for COP20-COP26.
Figure 4: Plot comparing the network von Neumann entropy of tweets for COP20-COP26.
Table 2: Pointwise estimates of network dimensionality and entropy based on the first 100 SVD values.
@@
20 14.0 0.9802473
21 7.0 0.97777313
22 2.0 0.97607696
23 3.0 0.97573924
24 9.0 0.9748669
25 3.0 0.9791759
26 2.0 0.9754415
Table 3: Pointwise estimates of network dimensionality and entropy based on the first 1000 SVD values.
@@
We found that the dimensionality of the network of Twitter users discussing the COP conference has been decreasing over time, though the decrease was not linear. Unexpectedly, the von Neumann entropy of the network did not decrease in this way, and instead it was at its highest in 2022 even though the dimensionality of the network was at its lowest.
@@
Interestingly, Falkenberg et al. expected to find polarisation during COP21, due to the signing of the Paris Agreement at COP21. Their Hartigan’s Dip Test for COP21 returned a significant result (p = 0.003), but they go on to claim that COP21 was not polarised despite this result. In our data, COP21 has lower dimensionality than the years before or after. It may be possible that the network as a whole became more polarised, which is captured by our data, but this effect had not yet occurred among the ”influencers” that Falkenberg et al selected. Our results support Falkenberg et al’s suggestion that the increase in polarisation they observed was due to an increase in the prominence of anti-climate and generally far-right influencers on Twitter, since COP26 was the conference with the lowest associated network dimensionality among Twitter users discussing it.
4.3 Simulated Data
Figure 5: Network dimensionality of the stochastic blockmodel as the link probabilities are changed.
Figure 6: von Neumann entropy of the stochastic blockmodel as the link probabilities are changed.
@@ 4.3 Simulated Data
As expected, increasing the chance of connection between the two blocks increases the dimensionality of the network (and therefore decreases the polarisation). The effect was consistent across all in-group link probabilities tested. This indicates that a potential social strategy to decrease polarsation could include facilitating the creation of connections between different groups.
@@
For cases where the in-group link probability was lower, the von Neumann entropy decreased as the out-group link probability increased. In cases where the in-group link probability was higher, the entropy remained consistent or increased as the out-group link probability increased. As such, von Neumann entropy may be a less reliable indicator of polarisation than $\hat{d}$.
Figure 7: Network dimensionality of the stochastic blockmodel as the sizes of the blocks are changed.
Figure 8: von Neumann entropy of the stochastic blockmodel as the sizes of the blocks are changed.
@@
We found that the dimensionality of the network increases slightly as one group becomes predominant in the network, but decreases strongly when one group is much larger than the other. This effect was consistent across all in-group link probabilities tested. The von Neumann entropy of the network also strongly decreased when one group was much larger than the other (99 to 1), but did not exhibit the same behaviour as the dimensionality when the group was only starting to become predominant (80 to 20, and 90 to 10). At low in-group link probabilities, the von Neumann entropy decreased as one group became predominant; at higher in-group link probabilities, the entropy either remained stable or increased slowly as one group became predominant.
@@
It is possible that our experiment did not decrease the group size far enough to trigger the effect expected by Esteban and Ray. However, it does demonstrate that polarisation does not linearly decrease as one group becomes predominant, as was expected, and that the behaviour of the stochastic block model is more complicated. This is also a useful result to be aware of because groups that are the target of political polarisation are often only a small percentage of the population, such as ethnic minority or refugee populations; as such, societies where this is the case may be structurally vulnerable to polarisation.
@@ 5 Conclusions
We have demonstrated a novel method for measuring polarisation through the embedded dimensionality of random dot-product graphs. This is a reliable and straightforward implementation of the correlation-based approach to polarisation suggested by Baldassarri and Gelman. Our method captured the presence of polarisation in all the scenarios where it was expected and had been found by other researchers, in both simulated data and real social media networks. The RDPG approach also allows us to easily see that the process of polarisation is occurring in a network, through its embedded dimensionality reducing, rather than relying on a binary test of whether the network was polarised or non-polarised.
@@
The von Neumann entropy of the network did not relate to the dimensionality as closely as we expected, though it did reflect major changes in the networks when they occurred. As such, we think it is best to use the embedded dimensionality of the network to measure its polarisation.
A major limitation of this method which could be improved is that it does not capture affective polarisation very well. Our method functionally considers any interaction between two nodes to be “good”; this means that it is not capable of capturing antagonistic interactions between nodes, and as a result it may overestimate the dimensionality of the network by mistaking brief antagonistic reactions for positive social bonds. There is a great deal of scope for integrating the concept of affective polarisation into our model, through methods such as using signed matrices and classification systems such as sentiment analysis to determine whether interactions in a social network are positive, negative, or neutral before determining the dimensionality.
Another possible extension of this method would be to implement a nonparametric two-sample hypothesis test[30], since this would allow a hypothesis test of whether the two networks are significantly different as additional evidence of polarisation having occurred. We believe that being able to observe the dimensionality of the graph alone is useful; however, we understand that sometimes a hypothesis test is demanded, and we believe this would help demonstrate that changes in the dimensionality of the network are significant.
Since we found some unexpected results when testing common axioms about polarisation, it would also be worthwhile to experiment more with the basics of the field using our method. For example, this paper has only explored the two-block case; many instances of online ”echo chambers” have a large number of groups who all hate people different to them, and it would be useful to see what happens to the dimensionality in such cases. Similarly, our testing on group prevalence showed a decrease in dimensionality when one group was 100 times the size of the other, but we did not test how the dimensionality changed as the size of the smaller group approaches zero.
5.0.1 Acknowledgements

### v3

@@ 2.3 Code
All scripts are available at https://doi.org/10.5281/zenodo.11043841. In Julia all packages can be installed by running Pkg.activate(“.”); Pkg.install().
3 Results
We apply our computational framework for polarisation to three different data sets: two from Twitter, and one consisting of simulated interaction networks. These have been chosen in order to show that our method works well, and to explore some common beliefs about polarisation in a new way.
@@ 3.1 Climate discussion in New Zealand Twitter
We obtained 12939 tweets by querying Twitter’s Academic API v2.0 for keywords related to climate change: climate, pollution, agw (anthropic global warming), CO2, and carbon. Some keywords used in other research, such as “COP 2x” for the international climate change conference, did not return many results when searched on New Zealand tweets. We restricted our search to original tweets geographically tagged as being from New Zealand, and tweets published after 2017, in order to investigate whether climate discussions in New Zealand were becoming more polarised.
We divided the data into two time windows that corresponded to equally sized, large networks: between 2017 and 2020, and 2020 to 2023. For each time frame, we built a network by considering each user (identified by their unique IDs) as a node, and any mention, reply, or quote tweet between two users as an edge. Retweets were excluded in order to maintain focus on the New Zealand network and prevent the network expanding to international discussions of climate change. There were 6767 tweets in the 2017-2020 network and 6172 in the 2020-2023 network; as such, the network sizes given are based on the number of nodes unless otherwise stated. We analysed the two networks independently.
We computed the results both in the original networks and in the original giant component of the network (the table of pointwise estimates below). We also computed results for 1000 bootstrapped samples of networks where we sampled (with repetition) the same amount of nodes of the original giant components of the graphs (the box graphs below). This bootstrapped sampling would allow to detect possible effect of sample limitations in the original graph – we did not find any notable sample limitations.
Figure 1: Plot comparing $\hat{d}$ of NZ climate change tweets in 2017-2020 to $\hat{d}$ in 2020-2023.
Figure 2: Plot comparing the SVD entropy of NZ climate change tweets in 2017-2020 to the entropy in 2020-2023.
@@
2017-2020 39 39 0.980229 0.979954
2020-2023 27 24 0.97439 0.97372
We find that $\hat{d}$ for the NZ Twitter discussion of climate change has decreased in the 2020-2023 network compared to the 2017-2020 network, indicating that the network was more polarised in the second time period than in the first. Similarly, the SVD entropy of the network is lower in the 2020-2023 period than in the 2017-2020 period. These results suggest that the complexity of the network decreased between these two time periods. This would support the argument that the political positions held by users are becoming narrower over time, which matches Baldassarri and Gelman’s definition of polarisation.
3.2 COP discussion in Twitter
Falkenberg et al. collected a very large corpus of tweets by querying Twitter’s Academic API v2.0 for tweets mentioning ”COP2x” where x was an integer between 0 and 6 (inclusive) [33]. Details about this data can be found in their original paper. They restricted their search to tweets in English, and covered the COP from 20 to 26 (years 2014 to 2022, with 2020 and 2021 skipped due to the Covid-19 pandemic). Their adjacency matrix was constructed based on whether a user $i$ retweeted tweets from a political influencer $j$. Their focus was whether there was noticeable division among users based on whether they were spreading true information about climate change or disinformation from climate change denialist influencers. To test for polarisation using our framework, we built a network for each year’s COP using the same data by considering each user (identified by their unique IDs) as a node, and any mention, reply, or quote tweet between two users as an edge. The network for each year was analysed independently. We then graphed $\hat{d}$ and the SVD entropy for each year in order to show how they, and therefore the polarisation of networks that discussed the COP conferences, changed over time.
@@ 3.2 COP discussion in Twitter
Figure 3: Plot comparing $\hat{d}$ of tweets for COP20-COP26.
Figure 4: Plot comparing the network SVD entropy of tweets for COP20-COP26.
Table 2: Pointwise estimates of $\hat{d}$ and entropy based on the first 100 SVD values.
@@
20 14 0.9802473
21 7 0.97777313
22 2 0.97607696
23 3 0.97573924
24 9 0.9748669
25 3 0.9791759
26 2 0.9754415
Table 3: Pointwise estimates of $\hat{d}$ and entropy based on the first 1000 SVD values.
@@
We found that $\hat{d}$ for the network of Twitter users discussing the COP conference has been decreasing over time, though the decrease was not linear. Unexpectedly, the SVD entropy of the network did not decrease in this way, and instead it was at its highest in 2022 even though $\hat{d}$ was at its lowest.
@@
Interestingly, Falkenberg et al. expected to find polarisation during COP21, due to the signing of the Paris Agreement at COP21 [33]. Their Hartigan’s Dip Test for COP21 returned a significant result (p = 0.003), but they go on to claim that COP21 was not polarised despite this result. In our data, COP21 has a lower $\hat{d}$ than the years before or after. It may be possible that the network as a whole became more polarised, which is captured by our data, but this effect had not yet occurred among the ”influencers” that Falkenberg et al selected. Our results support Falkenberg et al’s suggestion that the increase in polarisation they observed was due to an increase in the prominence of anti-climate and generally far-right influencers on Twitter, since COP26 was the conference with the lowest $\hat{d}$.
3.3 Synthetic Data
We wanted to explore what happens when common understandings of polarisation are evaluated using our approach. This both allows us to test these common understandings from new angles in order to see whether they hold up, and to find out whether there are specific cases where our methods may need to have corrections applied. We use stochastic block models for these experiments because they are straightforward to model using RDPGs while also being a good match for the experiments we wish to run. We focus on experimental computation of the statistical properties rather than analytical computation because the stochastic block models are the platform for our experiments rather than objects of interest in and of themselves.
3.3.1 Engagement Between Two Groups
The first common understanding we investigate is whether polarisation decreases as engagement between two groups increases (i.e. as echo chambers are removed). Understanding the effectiveness of this anti-polarisation strategy is important for implementing it in real-world social networks. We simulated a stochastic block network of 1000 nodes, split into two equally sized groups. We varied the probability of each node forming a connection within its group, and varied the probabilities of each node forming connections. In particular, we simulated a networks with in-group link probability of 0.3 to 0.45 with steps of 0.05. and between-group probabilities of 0.1, 0.05, and 0.01. Each combination of in-and between-group probabilities was simulated 100 times.
Figure 5: $\hat{d}$ of the stochastic blockmodel as the link probabilities are changed.
Figure 6: SVD entropy of the stochastic blockmodel as the link probabilities are changed.
@@ 3.3.1 Engagement Between Two Groups
As expected, increasing the chance of connection between the two blocks increases $\hat{d}$ (and therefore decreases the polarisation of the network). The effect was consistent across all in-group link probabilities tested. This indicates that a potential social strategy to decrease polarsation could include facilitating the creation of connections between different groups.
@@
For cases where the in-group link probability was lower, the SVD entropy decreased as the out-group link probability increased. In cases where the in-group link probability was higher, the entropy remained consistent or increased as the out-group link probability increased. As such, SVD entropy may be a less reliable indicator of polarisation than $\hat{d}$.
3.3.2 One Group Becoming Larger
The second common understanding we tested is that polarisation decreases as one of the groups becomes much bigger than the other. This has implications for real-world cases of polarisation where one group is much larger than the other, e.g. ethnic minorities or refugee communities.
We simulated a stochastic block network of 1000 nodes, split into two groups. We fixed probabilities of in-group linking between 0.3 and 0.45 with steps of 0.05, and fixed the probability of between-groups linking at 0.05. We varied the sizes of the two groups, progressively increasing the size of one group from 0.5 of the full network to 0.2, 0.1, and finally 0.01. Each combination of block size and in-group linking probabilities was simulated 100 times.
Figure 7: $\hat{d}$ of the stochastic blockmodel as the sizes of the blocks are changed.
Figure 8: SVD entropy of the stochastic blockmodel as the sizes of the blocks are changed.
@@ 3.3.2 One Group Becoming Larger
We found that $\hat{d}$ increases slightly as one group becomes predominant in the network, but decreases strongly when one group is much larger than the other. This effect was consistent across all in-group link probabilities tested. The SVD entropy of the network also strongly decreased when one group was much larger than the other (99 to 1), but did not exhibit the same behaviour as $\hat{d}$ when the group was only starting to become predominant (80 to 20, and 90 to 10). At low in-group link probabilities, the SVD entropy decreased as one group became predominant; at higher in-group link probabilities, the entropy either remained stable or increased slowly as one group became predominant.
It is important to note that this low $\hat{d}$ is because at the greatest group size difference, the smallest group only contains 10 nodes. This means that there is a risk of the smaller group becoming disconnected entirely from the larger group, and that even at its greatest likelihood of connecting the smaller group is only sparsely connected to the larger group. While it would be possible to redesign the experiment to focus on edge density rather than only connection probability, we think this would correspond conceptually to a change in how people in the network were making connections with each other, and we wished to keep this aspect the same in our experiment rather than varying it. We believe that the sparsity in our experiment is a useful feature, as some noted polarisations have one group that is much smaller than the other (ethnic minorities, refugee populations, and LGBT+ communities being examples); in these polarisations, effects created by one of the groups being very small are relevant, as the current understanding of polarisation would suggest that there is a mathematical upper limit to how severe polarisation against a small group can become – we suspect that this is not accurate.
@@
It is possible that our experiment did not decrease the group size far enough to trigger the effect expected by Esteban and Ray [30]. However, it does demonstrate that polarisation does not linearly decrease as one group becomes predominant, as was expected, and that the behaviour of the stochastic block model is more complicated.
@@ 4 Conclusions
We have demonstrated a novel method for measuring polarisation through the embedded dimensionality $\hat{d}$ of random dot-product graphs. This is a reliable and straightforward implementation of the correlation-based approach to polarisation suggested by Baldassarri and Gelman. Our method captured the presence of polarisation in all the scenarios where it was expected and had been found by other researchers, in both simulated data and real social media networks. In particular, the ability to observe the changes in $\hat{d}$ over time in data from the COP conferences provided a useful demonstration of how this method can be applied to longitudinal data. The RDPG approach also allows us to easily see that the process of polarisation is occurring in a network over time, through its embedded dimensionality reducing, rather than relying on a binary test of whether the network was polarised or non-polarised.
@@
The SVD entropy of the network did not relate to $\hat{d}$ as closely as we expected, though it did reflect major changes in the networks when they occurred. As such, we think it is best to use $\hat{d}$ of the network to measure its polarisation.
A major limitation of this method which could be improved is that it does not capture affective polarisation very well. Our method functionally considers any interaction between two nodes to be “good”; this means that it is not capable of capturing antagonistic interactions between nodes, and as a result it may overestimate $\hat{d}$ of the network by mistaking brief antagonistic reactions for positive social bonds. There is a great deal of scope for integrating the concept of affective polarisation into our model, through methods such as using signed matrices and classification systems such as sentiment analysis to determine whether interactions in a social network are positive, negative, or neutral before determining $\hat{d}$.
Another possible extension of this method would be to implement a nonparametric two-sample hypothesis test[47], since this would allow a hypothesis test of whether the two networks are significantly different as additional evidence of polarisation having occurred. We believe that being able to observe the embedded dimensionality of the graph alone is useful; however, we understand that sometimes a hypothesis test is demanded, and we believe this would help demonstrate that changes in the embedded dimensionality of the network are significant. In our tests using stochastic block models, we found an expected result that $\hat{d}$ increases as the probability of edges forming between the groups increases. This is evidence for the common belief about polarisation that connections between the groups decrease their polarisation. We also found an unexpected result that $\hat{d}$ does not straightforwardly increase when one group is much larger than the other, and in fact decreases greatly when one group is 100 times the size of the other. While this is somewhat a property of the size of the network and its sparsity, since the small group was very small and thus unable to make a large number of edges with the big group, we believe that it is an important case to have tested and that closer investigation of this scenario is warranted. We think that properties sparsity are important for modelling the real conditions of social networks at risk of polarisation, such as communities who have recently received refugees, and should be carefully considered in the model design to ensure that real effects are not being mathematically counteracted.
It would also be productive to experiment more with stochastic block models, since we only tested two possible scenarios. For example, this paper has only explored the two-block case; many instances of online ”echo chambers” have a large number of groups who all hate people different to them, and it would be useful to see what happens to the embedded dimensionality in such cases. Similarly, our testing on group prevalence showed a decrease in embedded dimensionality when one group was 100 times the size of the other, but we did not follow the smaller group’s size all the way to zero so we do not know whether the embedded dimensionality remains low or rebounds to a higher value.
4.1 Acknowledgements